OpenStudy (anonymous):

The function f has first derivative given by f'(x)=√x/1+x+x^3. What is the x-coordinate of the inflection point of the graph of f? A)1.008 B)0.473 C)0 D)-0.278 E)no point of inflection

OpenStudy (polaris_s0i):

is this you're equation? \[f^\prime(x) = \frac{\sqrt{x}}{1} + x +x^3\]

OpenStudy (anonymous):

to get the point of inflection, calculate the second derivative of the function and solve for it zeros, the value(s) of x that will make the function zero is the inflection point

OpenStudy (anonymous):

I got f''(x)=[1/2(√x)*(1+x+x^3)-(1+3x^2)*(√x)]/(1+x+x^3) but how do i continue this after setting it to 0?

OpenStudy (polaris_s0i):

oh this is you're function: \[f'(x)=\frac{\sqrt{x}}{1+x+x^3}\]

OpenStudy (polaris_s0i):

you need to square the bottom: \[f''(x)=\frac{\frac{1}{2}\sqrt{x}(1+x+x^3)-(1+3x^2)\sqrt{x}}{(1+x+x^3)^2}\] trying to figure out how to factor this beast.

OpenStudy (polaris_s0i):

sorry its this actually. we both didn't decrement g to -1/2 \[f''(x)=\frac{\frac{1}{2\sqrt{x}}(1+x+x^3)-(1+3x^2)\sqrt{x}}{(1+x+x^3)^2}\]